Five Years ago, Gemma was twice as oldas Tony is now. In seven years, Gemma will be five times as old as Tony is now. What are Gemma's and Tony current age's.
Answer by palanisamy(496)

Let the age of Gamma = x
Let the age of Tony = y
Given, Five Years ago, Gemma was twice as oldas Tony is now
x-5 = 2y ...(1)
Also. In seven years, Gemma will be five times as old as Tony is now.
x+7 = 5y ...(2)
(1)-(2)=>
-5-7 = 2y-5y
-12 = -3y
3y = 12
y = 12/3
y = 4
(1)=> x-5 = 2*4
x = 8+5 = 13
So, at present the age of Gamma = 13
and the age of Tony = 4